# Multilevel neural simulation-based inference

### v2

@@ B.6 Ablation study of the gradient adjustment technique
We further compare the training losses with and without gradient adjustment. With the gradient adjustment, the loss diverges, whereas with it, the loss curve exhibits convergence; see Figure˜10. To clarify the reason for this behaviour, we also provide an illustration of our gradient scaling procedure in Figure˜11.
We note that the optimisation requires some form of regularisation, and the one we used is one possible choice. In our experiments, the gradient adjustment approach performed the best among the options we tried (e.g., regularising the contribution of the difference term when it dominates the first term, or penalising high variance in the difference term). Moreover, it has the added benefit of not requiring any additional hyperparameters.

### v4

@@ B.6 Ablation study of the gradient adjustment technique
We further compare the training losses with and without gradient adjustment. With the gradient adjustment, the loss diverges, whereas with it, the loss curve exhibits convergence; see Figure˜11. To clarify the reason for this behaviour, we also provide an illustration of our gradient scaling procedure in Figure˜12.
We note that the optimisation requires some form of regularisation, and the one we used is one possible choice. In our experiments, the gradient adjustment approach performed the best among the options we tried (e.g., regularising the contribution of the difference term when it dominates the first term, or penalising high variance in the difference term).